\begin{thebibliography}{99}

\bibitem{ALWZ} R. Alweiss, S. Lovett, K. Wu and J. Zhang.
Improved bounds for the sunflower lemma.
STOC 2020: Proceedings of the 52nd Annual ACM SIGACT Symposium on Theory of Computing (2020) 624--630.

\bibitem{ERS}
P. Erd\H os. Problems and results in combinatorial analysis. (Italian summary) Colloquio Internazionale sulle Teorie Combinatorie (Rome, 1973), Tomo II, pp. 3–17. Atti dei Convegni Lincei, No. 17, Accad. Naz. Lincei, Rome, 1976.

\bibitem{EKR}
 P. Erd\H os, C. Ko and R. Rado. Intersection theorems for systems of finite sets. 
 {\it Quart. J. Math. Oxf.}
{\bf 2} (1961) 313--320.

\bibitem{ER}
P. Erd\H os and R. Rado.
Intersection theorems for systems of sets. 
{\it Journal of the London
Mathematical Society} {\bf 35} (1960) 85--90.


\bibitem{Fr3}
P. Frankl.
On intersecting families of finite sets.
{\it J. Combinatorial Theory Ser. A} {\bf 24} (1978) 146--161.

\bibitem{Fr-tau}
P. Frankl. On intersecting families of finite sets. {\it Bull. Austral. Math. Soc.} {\bf 21} (1980) 363--372.

\bibitem{Fr2}
P. Frankl.
Erd\H os-Ko-Rado theorem with conditions on the maximal degree.
{\it J. Combin. Theory Ser. A} {\bf 46} (1987) 252--263.

\bibitem{FR}
P. Frankl. Maximum degree and diversity in intersecting hypergraphs. {\it J. Combin. Theory Ser. B} {\bf 144} (2020) 81--94.

\bibitem{FrFu}  P. Frankl and Z. F\" uredi. A new generalization of the Erd\H os-Ko-Rado theorem.
{\it Combinatorica} {\bf 3}
(1983) 341--349.

\bibitem{FT2}
P. Frankl and N. Tokushige.
A note on Huang–Zhao theorem on intersecting families with large minimum degree.
{\it Discrete Math.}  {\bf 340} (2017) 1098--1103.

\bibitem{FT}
P. Frankl and N. Tokushige.
Invitation to intersection problems for finite sets. (English summary)
{\it J. Combin. Theory Ser. A} {\bf 144} (2016) 157--211.

\bibitem{FOT} P. Frankl, K. Ota and N. Tokushige.
Uniform intersecting families with covering number four.
{\it J. Combin. Theory Ser. A} {\bf 71} (1995) 127--145.

\bibitem{Fu}
Z. F\"uredi.
Erd\H os-Ko-Rado type theorems with upper bounds on the maximum degree. Algebraic methods in graph theory, Vol. I, II (Szeged, 1978), pp. 177–207,
Colloq. Math. Soc. János Bolyai, 25, North-Holland, Amsterdam-New York, 1981.

\bibitem{Fu2}
Z. F\"uredi.
An intersection problem with 6 extremes.
{\it Acta Math. Hungar.} {\bf 42} (1983) 177--187.

\bibitem{Fu3}
Z. F\" uredi. On finite set-systems whose every intersection is a Kernel of a star. {\it Discrete Math.}
{\bf 47} (1983) 129--132.

\bibitem{HZ} H. Huang and Y. Zhao. Degree versions of the Erd\H os–Ko–Rado theorem and Erd\H os hypergraph matching conjecture. {\it J. Combin. Theory Ser. A} 
{\bf 150} (2017) 233--247.

\bibitem{KM} P. Keevash and D. Mubayi. Stability theorems for cancellative hypergraphs.
{\it J. Combin. Theory Ser. B} {\bf 92} (2004) 163--175.

\bibitem{KMV} A. Kostochka, D. Mubayi and J. Verstra\"ete.
Tur\'an Problems and Shadows I: Paths and Cycles.
{\it J. Combin. Theory Ser. A} {\bf 129} (2015) 57--79.

\bibitem{Ku1} A. Kupavskii. Degree versions of theorems on intersecting families
via stability. {\it J.~Combin. Theory Ser. A}
{\bf 168} (2019) 272--287.

\bibitem{KU} A. Kupavskii. Diversity of uniform intersecting families. {\it European J. Combin.} {\bf 74} (2018) 39--47.

\bibitem{LP} N. Lemons and C. Palmer.
The unbalance of set systems. 
{\it Graphs Combin.} {\bf 24} (2008) 361--365.

\bibitem{MZ} D. Mubayi and Y. Zhao. Co-degree density of hypergraphs.
{\it J. Combin. Theory Ser. A} {\bf 114} (2007) 1118--1132.

\bibitem{MZ2} D. Mubayi and Y. Zhao. Forbidding complete hypergraphs as traces. {\it Graphs  Combin.} {\bf 23} (2007) 667--679.

\end{thebibliography}

\documentclass[12pt]{article}
\usepackage[a4paper]{geometry}
\usepackage{latexsym,amssymb,upref,amsmath,amsthm, amsfonts,authblk}
\usepackage{color}
\usepackage{fullpage}
\usepackage{url}

\newtheorem{thm}{Theorem} 
\newtheorem{lemma}[thm]{Lemma}
 
\newtheorem{defn}[thm]{Definition}
\newtheorem{claim}[thm]{Claim}
\newtheorem{prop}[thm]{Proposition}

\newtheorem*{remark}{Remark}



\newcommand{\A}{\mathcal{A}}
\newcommand{\B}{\mathcal{B}}

\newcommand{\h}{\mathcal{H}}

\title{
	Maximum size intersecting families of bounded minimum positive co-degree
}

\begin{document}

\author{
J\'ozsef Balogh\thanks{Department of Mathematics, University of Illinois at Urbana-Champaign, IL, 
USA, and MIPT, Russian Federation. Email: \texttt{jobal@illinois.edu}.
Partially supported by NSF Grant DMS-1764123 and Arnold O. Beckman Research Award (UIUC) Campus Research Board 18132, Simons Fellowship 
and the Langan Scholar Fund (UIUC).}
\qquad
Nathan Lemons\thanks{Theoretical Division, Los Alamos National Laboratory, Email: \texttt{nlemons@lanl.gov}.}
\qquad
Cory Palmer\thanks{Department of Mathematical Sciences, University of Montana. Email: \texttt{cory.palmer@umontana.edu}.
Research supported by a grant from the Simons Foundation \#712036.}

}
\date{
\today}
\maketitle

\begin{abstract}
 Let $\mathcal{H}$ be an $r$-uniform hypergraph. The \emph{minimum positive co-degree} of $\mathcal{H}$, denoted by $\delta_{r-1}^+(\mathcal{H})$, is the minimum $k$ such that if $S$ is an $(r-1)$-set contained in a hyperedge of $\mathcal{H}$, then $S$ is contained in at least $k$ hyperedges of $\mathcal{H}$. For $r\geq k$ fixed and $n$ sufficiently large, we determine the maximum possible size of an intersecting $r$-uniform $n$-vertex hypergraph with minimum positive co-degree $\delta_{r-1}^+(\mathcal{H}) \geq k$ and characterize the unique hypergraph attaining this maximum. This generalizes the Erd\H os-Ko-Rado theorem which corresponds to the case $k=1$.  Our proof is based on the delta-system method. 
\end{abstract}

\section{Introduction}

A hypergraph $\mathcal{H}$ is {\it intersecting} if for every pair of hyperedges $h,h' \in E(\mathcal{H})$ we have $h \cap h' \neq \emptyset$. The celebrated theorem of Erd\H os, Ko and Rado~\cite{EKR} gives that for $n \geq 2r$, the maximum size of an intersecting $r$-uniform $n$-vertex hypergraph is $\binom{n-1}{r-1}$. 
The Erd\H os-Ko-Rado theorem is a cornerstone of extremal combinatorics and has many proofs, extensions and generalizations, see the excellent survey of Frankl and Tokushige~\cite{FT} for a history of extremal problems for intersecting hypergraphs.
We call the unique hypergraph achieving the maximum in the Erd\H os-Ko-Rado theorem a {\it maximal star}, i.e., the hypergraph of all hyperedges containing a given vertex.

The {\it degree} of a set of vertices $S$ in a hypergraph $\mathcal{H}$ is the number of hyperedges containing $S$, i.e., $|\{ h \in E(\mathcal{H}) : S \subseteq h\}|$. 
Denote by $\delta_s(\h)$ the minimum degree of an $s$-element subset of the vertices of $\h$. In this way, $\delta_1(\h)$ is the standard minimum degree of a vertex in $\h$.

Huang and Zhao~\cite{HZ} considered a minimum degree version of the Erd\H os-Ko-Rado theorem. In particular, they proved that for $n \geq 2r+1$, if $\mathcal{H}$ is an intersecting $r$-uniform $n$-vertex hypergraph, then $\mathcal{H}$ has minimum degree  $\delta_1(\h) \leq \binom{n-2}{r-2}$. The Huang-Zhao~\cite{HZ} proof uses the linear algebra method and later a combinatorial proof was given by Frankl and Tokushige~\cite{FT2} for $n \geq 3r$.
Kupavskii~\cite{Ku1} gave an extension of this result and showed that for $t <r$ and $n \geq 2r+3t/(1-t/r)$, every intersecting $r$-uniform $n$-vertex hypergraph $\mathcal{H}$ satisfies $\delta_t(\mathcal{H}) \leq \binom{n-t-1}{r-t-1}$. 

In the more general hypergraph setting, Mubayi and Zhao~\cite{MZ} introduced the notion of co-degree Tur\'an numbers, i.e., the maximum possible value of $\delta_{r-1}(\mathcal{H})$ among all $r$-uniform $n$-vertex hypergraphs $\h$ not containing a specified subhypergraph $\mathcal{F}$. In their paper they give several results that show that the co-degree extremal problem behaves differently from the classical Tur\'an problem.

Motivated by the degree versions of the Erd\H os-Ko-Rado theorem and co-degree Tur\'an numbers we propose studying the following  hypergraph degree condition.

\begin{defn}
    Let $\mathcal{H}$ be a non-empty $r$-uniform hypergraph. The \emph{minimum positive co-degree} of $\mathcal{H}$, denoted $\delta_{r-1}^+(\mathcal{H})$, is the maximum $k$ such that if $S$ is an $(r-1)$-set contained in a hyperedge of $\mathcal{H}$, then $S$ is contained in at least $k$ distinct hyperedges of $\mathcal{H}$. 
\end{defn}

Note that the empty hypergraph is a degenerate case; for simplicity we define its positive co-degree to be zero.

As an example, let us examine hypergraphs that contain no $F_5 = \{abc,abd,cde\}$ to compare the co-degree and positive co-degree settings. Frankl and F\" uredi~\cite{FrFu} (see~\cite{KM} for a strengthening) showed that the complete balanced tripartite $3$-uniform hypergraph has the maximum number of hyperedges among all $3$-uniform $n$-vertex $F_5$-free hypergraphs, for $n$ sufficiently large. This construction has minimum co-degree $0$ and it is easy to see that minimum co-degree at least $2$ guarantees the existence of an $F_5$. On the other hand, the balanced tripartite hypergraph is $F_5$-free and has minimum positive co-degree $n/3$ and it can be shown that minimum positive co-degree strictly greater than $n/3$ implies the existence of an $F_5$.

Note that for ordinary graphs (i.e.\ 2-uniform hypergraphs), the minimum positive co-degree is simply the minimum degree of the non-isolated vertices, which in many extremal problems we may assume is equal to the minimum degree. This suggests positive co-degree as a reasonable notion of ``minimum degree'' in a hypergraph.

The positive co-degree condition has appeared in several other contexts. For example, in~\cite{KMV} the term {\it $d$-full} was used and the authors gave some simple lemmas for hypergraphs with minimum positive co-degree (in the course of proving theorems about extremal numbers for hypergraphs).

In this paper we investigate the maximum size of an intersecting $r$-uniform $n$-vertex hypergraph with positive co-degree at least $k$. As the condition $\delta_{r-1}^+(\mathcal{H}) \geq 1$ is vacuous, the maximum in this case is $\binom{n-1}{r-1}$ as given by the Erd\H os-Ko-Rado theorem.  The unique construction achieving this bound has minimum positive co-degree 1.  On the other hand, as shown in Proposition~\ref{codegree-uniformity-bound}, in an intersecting hypergraph the uniformity gives an upper bound on the minimum positive co-degree, i.e., $r \ge k$. Thus the range of interest for our problem is $2\leq k \leq r$.  In this range we prove that for $n$ sufficiently large the maximum-size intersecting hypergraph with minimum positive co-degree $k$ is given by the following hypergraph.

\begin{defn}
Given integers $r \geq k\geq 1$ an \emph{($r$-uniform) $k$-kernel system} is a hypergraph $\mathcal{H}$ on vertex set $V$ with edges $\mathcal{E}=\{E\in \binom{V}{r}: |E\cap X|\geq k\}$, were $X$ is a distinguished subset of $V$ of size $2k-1$.  The set $X$ is called the \emph{kernel} of $\mathcal{H}$.
\end{defn}

Clearly a $k$-kernel system is intersecting.
Observe that the number of hyperedges in an $r$-uniform $n$-vertex $k$-kernel system $\h$ is 
\[
|E(\h)| = \sum_{i=k}^{\max\{r, 2k-1\}} \binom{2k-1}{i} \binom{n-2k+1}{r-i} \geq \binom{2k-1}{k} \binom{n-2k+1}{r-k}= \Omega(n^{r-k}).
\]

Note that a $1$-kernel system is the hypergraph consisting of all hyperedges containing a fixed vertex $x$, i.e., the maximal hypergraph in the Erd\H os-Ko-Rado theorem. Interestingly, $k$-kernel systems appear as solutions to  maximum degree versions of the Erd\H os-Ko-Rado theorem. Let us give three examples. 

First, a special case of a more general theorem of Frankl~\cite{Fr2} implies that if $\mathcal{H}$ is a maximum-size intersecting $r$-uniform $n$-vertex hypergraph with maximum degree at most $2\binom{n-3}{r-2} + \binom{n-3}{r-3}$, then $\h$ is a $2$-kernel system, provided $n$ is large enough. 

Second, Erd\H os, Rothschild and Szemer\'edi (see~\cite{ERS}) posed the following problem: determine the maximum size of an intersecting $r$-uniform $n$-vertex hypergraph $\mathcal{H}$ such that each vertex contained in at most $c|E(\mathcal{H})|$ hyperedges for $r \geq 3$ and $0 < c <1$. They proved when $c=2/3$ and $n$ large, then a $2$-kernel system is the unique hypergraph attaining this maximum. Frankl~\cite{Fr3} showed that for $2/3 \leq c <1$ and $n$ large enough, $\h$ has no more hyperedges than a $2$-kernel system. 
For $3/5< c < 2/3$ and $n$ large enough, F\"uredi~\cite{Fr3} showed that a $3$-kernel system is one of six non-isomorphic hypergraphs attaining this maximum. 
In the case when $1/2< c \leq 3/5$ and $n$ large enough, Frankl~\cite{Fr3} showed that $\h$ has no more hyperedges than a $3$-kernel system, although the unique hypergraph attaining this maximum is not isomorphic to a $3$-kernel system.

Third, Lemons and Palmer~\cite{LP} proved that $3$-kernel systems are the $r$-uniform $n$-vertex hypergraphs with the largest {\it diversity}, i.e., the difference between the number of hyperedges and the maximum degree
for $n$ large enough (see~\cite{FR,KU} for improvements to the threshold on $n$).

The main result of our paper is as follows:

 \begin{thm}\label{general-k-thm}
    Let $\mathcal{H}$ be an intersecting $r$-uniform $n$-vertex hypergraph with minimum positive co-degree $\delta_{r-1}^+(\mathcal{H}) \geq k$ where $1 \leq k \leq r$. If $\mathcal{H}$ has the maximum number of hyperedges, then $\mathcal{H}$ is a $k$-kernel system for $n$ sufficiently large.
\end{thm}
Theorem~\ref{general-k-thm} holds for $n$ large, roughly double exponential in $r$.
In Section~\ref{small-n-section} we give two results that suggest that Theorem~\ref{general-k-thm} should hold for $n$ at least $cr^{k+2}$, where $c$ is a polynomial in $k$.
It would be interesting to further refine the range of $n$ as a function of $r$ and $k$ where our results hold. Also, we only considered the positive co-degree of $(r-1)$-sets. We can define $\delta^+_s(\mathcal{H})$ to be the minimum $k$ such that if $S$ is an $s$-set contained in a hyperedge of $\mathcal{H}$, then $S$ is contained in at least $k$ distinct hyperedges. There may be interesting problems to be considered under this more general condition. 

\section{Proof of Theorem~\ref{general-k-thm}}

First, let us observe that the uniformity of an intersecting hypergraph is always at least the minimum positive co-degree.

\begin{prop}\label{codegree-uniformity-bound}
    If $\mathcal{H}$ is a non-empty intersecting $r$-uniform $n$-vertex hypergraph with minimum positive co-degree $\delta^+_{r-1}(\mathcal{H}) \geq k$, then $r \geq k$.
\end{prop}

\begin{proof}
Assume, for the sake of a contradiction, that $k > r$.
    Let $h = \{x_1,x_2,\dots, x_r\}$ be a hyperedge of $\mathcal{H}$. The $(r-1)$-set $h \setminus \{x_1\}$ has co-degree at least $k$, so there is a vertex $x_{r+1} \not \in h$ such that $\left(h \setminus \{x_1\}\right) \cup \{x_{r+1}\} $ is a hyperedge of $\mathcal{H}$. Similarly, the $(r-1)$-set $\left(h \setminus \{x_1,x_2\}\right) \cup \{x_{r+1}\}$ has co-degree at least $k$, so there is a vertex $x_{r+2} \not \in h\cup \{x_{r+1}\}$ such that  $\left(h \setminus \{x_1,x_2\}\right) \cup \{x_{r+1},x_{r+2}\}$ is a hyperedge of $\mathcal{H}$. Because $k > r$, we can repeat this process to obtain a hyperedge $\left(h \setminus \{x_1,\dots, x_r\}\right) \cup \{x_{r+1},\dots, x_{2r}\} = \{x_{r+1},\dots, x_{2r}\}$ that is in $\mathcal{H}$. Now we have disjoint hyperedges $h$ and $\{x_{r+1},\dots, x_{2r}\}$ in $\h$ which contradicts the intersecting property.
\end{proof}

An $r$-uniform hypergraph $\mathcal{S}$ is a {\it sunflower} if every pairwise intersection of the hyperedges is the same set $Y$, called the {\it core} of the sunflower. We call the sets $h \setminus Y$ for $h \in E(\mathcal{S})$ the {\it petals} of the sunflower $\mathcal{S}$. Note that the petals are pairwise disjoint. Denote the size of the core of a sunflower $\mathcal{S}$ by $c(\mathcal{S})$.


Let $f(r,p)$ denote the minimum integer such that an $r$-uniform hypergraph with $f(r,p)$ hyperedges contains a sunflower with $p$ petals. 
The Sunflower Lemma of Erd\H os and Rado~\cite{ER} claims that $f(r,p) \leq r! (p-1)^r$.
The determination of $f(r,p)$ is a well-known open problem in combinatorics. A recent breakthrough by Alweiss, Lovett, Wu and Zhang~\cite{ALWZ} gives a bound on $f(r,p)$ of about $(\log r)^{r(1+o(1))}$.

In general we cannot force a sunflower to have a core of a specified size unless we increase the number of hyperedges in the host hypergraph. Mubayi and Zhao (Lemma~6 in~\cite{MZ2}) gives conditions for the existence of a sunflower with a core of bounded size.

\begin{lemma}[Mubayi and Zhao,~\cite{MZ2}]\label{k-core}
Fix integers $r\geq 3$, $k\geq 1$ and $p \geq 1$ and let $C=C(r,p)$ be a large enough constant.
If $\mathcal{G}$ is an $r$-uniform $n$-vertex hypergraph with 
\[
|E(\mathcal{G})| \geq C n^{r-k-1},
\]
then  $\mathcal{G}$ contains a sunflower with $p$ petals and core of size at most $k$.  
\end{lemma}

Observe that Lemma~\ref{k-core} is sharp in the order of magnitude of $n$. Indeed, the $r$-uniform $n$-vertex hypergraph consisting of all hyperedges containing a fixed set $Y$ of $k+1$ vertices contains $\binom{n-k-1}{r-k-1}$ hyperedges, but no sunflower with a core of size at most $k$ as any two hyperedges intersect in at least $k+1$ vertices.
We remark that the problem to determine the best constant $C$ in Lemma~\ref{k-core} is interesting in its own right. In the Appendix at the end of the paper we give a new proof of Lemma~\ref{k-core} that gives an improvement to $C$.

We will need a lower bound on the size of a core of a sunflower in an intersecting hypergraph.

\begin{lemma}\label{sunflower-transversal}
If $\mathcal{S}$ is a sunflower with at least $r+1$ petals in an intersecting $r$-uniform hypergraph $\mathcal{G}$ with $\delta_{r-1}^+(\mathcal{G})\geq k$, then the core $Y$ of $\mathcal{S}$ satisfies $|Y|\geq k$.
\end{lemma}

\begin{proof}
For the sake of contradiction, assume that the core $Y$ of $\mathcal{S}$ is small, i.e., $|Y|<k$.
Observe that $Y$ is a transversal of $\mathcal{G}$, i.e., every hyperedge of $\mathcal{G}$ intersects $Y$. Indeed, 
as the petals of the sunflower $\mathcal{S}$ are pairwise vertex-disjoint, each hyperedge of $\mathcal{G}$ must intersect the core $Y$ in order to intersect each of the at least $r+1$ hyperedges associated with the petals of the sunflower.

Now let $Y'$ be a minimum transversal in $\mathcal{G}$. Thus $|Y'| \leq |Y| < k$ and the minimality of $Y'$ guarantees the existence of a hyperedge $h$ that intersects $Y'$ in exactly one element.
 The $(r-1)$-set $h \setminus Y'$ is contained in at most $k-1$ hyperedges of $\mathcal{G}$; one for each element of $Y'$. This contradicts the positive co-degree condition on $\mathcal{G}$.
\end{proof}


\begin{proof}[Proof of Theorem~\ref{general-k-thm}.]
Let $\mathcal{H}$ be an intersecting $r$-uniform $n$-vertex hypergraph with minimum positive co-degree $\delta_{r-1}^+(\mathcal{H}) \geq k$ where $1 \leq k \leq r$. Moreover, suppose that $\h$ has the maximum number of hyperedges. We will show that $\h$ is a $k$-kernel system for $n$ sufficiently large.

We have observed that a $k$-kernel system has minimum positive co-degree at least $k$, so we may assume that
\[
|E(\mathcal{H})| \geq \binom{2k-1}{k} \binom{n-2k+1}{r-k} = \Omega(n^{r-k}).
\]
Therefore, for $n$ large enough, Lemmas~\ref{k-core} and~\ref{sunflower-transversal} guarantees the existence of a sunflower $\mathcal{S}$ with 
$p=(r+1)  r^{k-1}$ petals and core of size $k$. 
Denote the core of $\mathcal{S}$ by $Y=\{y_1,y_2,\dots, y_k\}$.

Note that in order to apply Lemma~\ref{k-core} we need that the following inequality is satisfied:
\[
\binom{2k-1}{k}\binom{n-2k+1}{r-k} \geq Cn^{r-k-1},
\]
where $C=C(r,p)$ is the constant from Lemma~\ref{k-core}. This is satisfied when
\[
n \geq  \frac{(2r-2k)^{r-k}}{\binom{2k-1}{k}} C. 
\]
The value $C = (pr2^r)^{2^r}$ given in \cite{MZ2} follows from a 
theorem of F\"uredi \cite{Fu3}. 


\begin{claim}\label{sunflower-pair}
There is a set of vertices $Z = \{z_1,z_2,\dots, z_{k-1}\}$ such that $Z \cap Y = \emptyset$ and $Z \cup \{y_k\}$ is the core of a sunflower with $r+1$ petals.
\end{claim}

\begin{proof}
We will prove the following stronger claim: For $0 \leq i \leq k-1$, there is a set of vertices $Z_i=\{z_1,z_2,\dots, z_{i}\}$ such that $Y \cap Z_i = \emptyset$ and $Z_i \cup \{y_k,y_{k-1},\dots, y_{i+1}\}$ is the core of a sunflower $\mathcal{S}_i$ with $(r+1)  r^{k-1-i}$ petals. The claim follows from the case $i=k-1$.

We proceed by induction on $i$. The base case $i =0$ is immediate as $Z_0=\emptyset$ and $\mathcal{S}_0 = \mathcal{S}$ is a sunflower with core $Z_0 \cup \{y_k,y_{k-1},\dots, y_1\} = Y$ with $(r+1)  r^{k-1}$ petals. 
Now suppose $i>0$ and the statement holds for $i-1$. Let $\mathcal{S}_{i-1}$ be a sunflower given by the inductive hypothesis.

For each petal $P$ in $\mathcal{S}_{i-1}$ consider the $(r-1)$-set $P\cup Z_{i-1} \cup \{y_k,\dots, y_{i+1}\} = P\cup Z_{i-1} \cup \{y_k,\dots, y_i\} \setminus \{y_{i}\}$. By the positive co-degree condition on $\mathcal{H}$, the set $P\cup Z_{i-1} \cup \{y_k,\dots, y_{i+1}\}$ is contained in $k$ hyperedges of $\mathcal{H}$. Therefore, as $i \leq k-1$, there is a vertex $x(P)$ such that $x(P) \not \in \{y_1,y_2,\dots, y_i\}$ and $\{x(P)\} \cup P\cup Z_{i-1} \cup \{y_k,\dots, y_{i+1}\}$ is a hyperedge of $\mathcal{H}$.

Now suppose there are distinct vertices $x_1,x_2,\dots, x_{r+1}$ among the vertices in $\{x(P) : P \text{ is a petal in } \mathcal{S}\}$. Let $P_1,P_2,\dots, P_{r+1}$ be the petals corresponding to these vertices, i.e., $\{x_j\} \cup P_j \cup Z_{i-1} \cup \{y_k,\dots, y_{i+1}\} \in E(\mathcal{H})$ for $j=1,2,\dots, r+1$. Then $Z_{i-1} \cup \{y_k,\dots, y_{i+1}\}$ is the core of size $k-1$ of a sunflower with petals $P_j \cup \{x_j\}$ for $j=1,2,\dots, r+1$ in $\mathcal{H}$. This contradicts Lemma~\ref{sunflower-transversal}. Therefore, there are at most $r$ distinct vertices among the vertices in $\{x(P) : P \text{ is a petal in } \mathcal{S}_{i-1}\}$. This implies that there is a vertex $x$ that is the vertex $x(P)$ for at least $\frac{1}{r}|E(\mathcal{S}_{i-1})| \geq (r+1)  r^{k-2-(i-1)}$ petals $P$ in $\mathcal{S}_{i-1}$. Put $z_i=x$ and $Z_i = \{z_1,z_2,\dots, z_i\}$ and let $\mathcal{S}_i$ be the sunflower consisting of $(r+1)  r^{k-1-i}$ hyperedges of $\mathcal{S}_{i-1}$ containing $x = z_i$. Observe that $Z_i \cup \{y_k,\dots, y_{i+1}\}$ is the core of the sunflower $\mathcal{S}_i$ with $(r+1)  r^{k-1-i}$ petals.
\end{proof}

Let $\mathcal{S}_Z$ be a sunflower with $r+1$ petals and core $Z \cup \{y_k\}$ given by Claim~\ref{sunflower-pair}. 
There are at most $(r+1)(r-k)+(k-1)$ vertices
disjoint from $Y$ spanned by $\mathcal{S}_Z$.
As $\mathcal{S}$ has $(r+1) r^{k-1}$ petals, we may choose $r+1$ petals of $\mathcal{S}$ that are vertex-disjoint from the vertices of $\mathcal{S}_Z$. Call the resulting sunflower $\mathcal{S}_Y$. Note that $\mathcal{S}_Y$ has $r+1$ petals and core $Y$.

\begin{claim}\label{kernel-claim}
For every petal $P$ in $\mathcal{S}_Z$ and every $y \in Y$ we have that $P \cup Z \cup \{y\}$ is a hyperedge in $\mathcal{H}$.
\end{claim}

\begin{proof}
Observe that the $(r-1)$-set $P \cup Z$ is contained in the hyperedge $P \cup Z \cup \{y_k\}$, so by the positive co-degree condition $P \cup Z$ is contained
in $k$ hyperedges of $\mathcal{H}$. Moreover, each of these hyperedges must intersect every hyperedge in the sunflower $\mathcal{S}_Y$. As $\mathcal{S}_Y$ has at least $2$ petals, each of the $k$ hyperedges containing $P \cup Z$ must contain a distinct vertex of $Y$.
\end{proof}

We now continue with a technical claim that will imply the theorem.

\begin{claim}\label{technical-claim} 
For every $k$-set $T \subset Y\cup Z$ we have:
\begin{enumerate}
\item[(1)] $Q \cup T \in E(\mathcal{H})$ for every petal $Q$ of $\mathcal{S}_Y$,

\item[(2)] $((Y \cup Z)\setminus T) \cup \{s\} \cup P  \in E(\mathcal{H})$ for every $s \in T$ and petal $P$ of $\mathcal{S}_Z$.
\end{enumerate}
\end{claim}

\begin{proof}
We proceed by induction on $t=|T \cap Z|$. Note that $t \leq k-1$.
When $t=0$ we have that $T =Y$, then (1) is immediate as $Q \cup Y \in E(\mathcal{S}_Y) \subset \mathcal{H}$ and (2) follows from Claim~\ref{kernel-claim}. 

Let $t>0$ and suppose the statement of the claim holds for all smaller values of $t$. As $0<  t\leq k-1$, there exists a $z \in Z \cap T$ and a $y \in Y\setminus T$. Fix an arbitrary petal $Q$ of $\mathcal{S}_Y$. 
Put $T' =  T \cup \{y\} \setminus \{z\}$ and note that $|T' \cap Z| = t-1$. Therefore, by induction, we have $Q \cup T' \in E(\mathcal{H})$ and $((Y\cup Z) \setminus T') \cup \{s'\} \cup P \in E(\mathcal{H})$ for every $s' \in T'$ and petal $P$ of $\mathcal{S}_Z$.

By the positive co-degree condition, the $(r-1)$-set $Q \cup T' \setminus \{y\}$ is contained in at least $k$ hyperedges. Moreover, $Q \cup T' \setminus \{y\}$ is disjoint from the hyperedges of the form $((Y \cup Z) \setminus T') \cup \{y\} \cup P$ where $P$ is a petal of $\mathcal{S}_Z$.
As $\mathcal{S}_Z$ has $r+1$ petals and $\h$ is intersecting, this implies that
the $k$ hyperedges containing $Q \cup T' \setminus \{y\}$ each intersect the $k$-set $((Y \cup Z) \setminus T) \cup \{y\}$. In particular, $(Q \cup T' \setminus \{y\}) \cup \{z\} = Q \cup T$ is a hyperedge of $\mathcal{H}$. This proves (1). 

In order to prove (2), let us fix an arbitrary petal $P$ of $\mathcal{S}_Z$.
Observe that the $(r-1)$-set 
\[
((Y \cup Z) \setminus T) \cup P =((Y \cup Z) \setminus ( T' \cup \{z\} \setminus \{y\})) \cup P  = ((Y \cup Z) \setminus T') \setminus \{z\} \cup \{y\} \cup P
\]
is contained in 
the hyperedge $(Y \cup Z) \setminus T' \cup \{y\} \cup P \in E(\mathcal{H})$ whose existence is given by the inductive hypothesis on (2) with $y=s' \in T'$. Therefore, the positive co-degree condition guarantees that the $(r-1)$-set 
$((Y \cup Z) \setminus T) \cup P$ is contained in $k$ hyperedges. In order for these hyperedges to intersect the $r+1$ hyperedges $Q \cup T$ for each petal $Q$ of $\mathcal{S}_Y$, we have that each set of the form $((Y \cup Z) \setminus T) \cup \{s\} \cup P$ for $s \in T$ must be a hyperedge of $\mathcal{H}$.
\end{proof}

We are now ready to complete the proof of Theorem~\ref{general-k-thm}. Suppose that there is a hyperedge $h \in E(\mathcal{H})$ such that $|h \cap (Y \cup Z)| \leq k-1$. Then there exists a $k$-set $T \subset Y \cup Z$ such that $T$ is disjoint from $h$. Moreover, as $\mathcal{S}_Y$ has at least $r+1$ petals, there is a petal $Q$ in $\mathcal{S}_Y$ that is disjoint from $h$. By Claim~\ref{technical-claim} we have that $T \cup Q  \in E(\mathcal{H})$ which is disjoint from $h \in E(\mathcal{H})$. This violates the intersecting property of $\mathcal{H}$, a contradiction.

Therefore, every hyperedge $h\in E(\mathcal{H})$ intersects $Y \cup Z$ in at least $k$ vertices. This implies that $\mathcal{H}$ is a subhypergraph of a $k$-kernel system, i.e., as $\mathcal{H}$ is edge-maximal, it is exactly a $k$-kernel system.
\end{proof}

\begin{remark} \normalfont
Observe that the proof of Theorem~\ref{general-k-thm} gives a stability result. In particular, if $\mathcal{H}$ has enough edges to apply Lemma~\ref{k-core}, then we have that $\mathcal{H}$ is a subhypergraph of a $k$-kernel system.
\end{remark}

\section{Improved thresholds on $n$}\label{small-n-section}

We now show that in the case $k\leq 3$, Theorem~\ref{general-k-thm} holds for $n \geq c r^{k+2}$. In Theorem~\ref{general-k-thm} we need $n$ to be at least double exponential in $r$.
Recall that two hypergraphs $\mathcal{A}$ and $\mathcal{B}$ are {\it cross-intersecting} if for every pair of hyperedges $A \in E(\mathcal{A})$ and $B \in E(\mathcal{B})$ we have $A \cap B \neq \emptyset$.
Also, a {\it transversal} for a hypergraph $\mathcal{H}$ is a set of vertices $T$ such that $T \cap h \neq \emptyset$ for every hyperedge $h \in E(\h)$. The {\it transversal number} $\tau(\h)$ is the minimum $t$ such that there is a transversal $T$ of $\h$ of size $t$.

We begin with a simple bound on the size of an intersecting hypergraph $\mathcal{H}$ with transversal number $\tau(\mathcal{H})=t$. Stronger results for $\tau(\mathcal{H})=3$ and $\tau(\mathcal{H})=4$ are given by Frankl~\cite{Fr-tau} and Frankl, Ota and Tokushige~\cite{FOT}, but we include an argument for the sake of  completeness and as our argument holds for all $n$ and $t$.

\begin{lemma}\label{tau-bound}
    Fix $n \geq r \geq t$.
	Let $\h$ be an intersecting $r$-uniform $n$-vertex hypergraph with transversal number $\tau(\h) \geq t$. Then
	\[
	|E(\h)| \leq r^{t} \binom{n-t}{r-t}.	
	\]
\end{lemma}

\begin{proof}
    Let us construct a $t$-uniform hypergraph $\mathcal{T}$ with $|E(\mathcal{T})| \leq r^t$ such that for every $h \in E(\mathcal{H})$ there exists a $h' \in E(\mathcal{T})$ with $h' \subset h$. The existence of $\mathcal{T}$ immediately implies the lemma as $|E(\h)| \leq |E(\mathcal{T})| \binom{n-t}{r-t}$.
    
    We proceed iteratively. First select an arbitrary hyperedge $h_1 \in E(\h)$. For each vertex $v_1 \in h_1$, the set $\{v_1\}$ is not a transversal of $\h$, so there is a hyperedge $h_2 \in E(\h)$ that is disjoint from $\{v_1\}$.
    For each vertex $v_2 \in h_2$, the set $\{v_1,v_2\}$ is not a transversal of $\h$, so there is a hyperedge $h_3 \in E(\h)$ that is disjoint from $\{v_1,v_2\}$. We continue this process to select a set of $t$ distinct vertices $v_1,v_2,\dots, v_t$. Let $\mathcal{T}$ be the collection of all $t$-sets constructed in this way. 
    Note that in each step there are at most $r$ choices for the vertex $v_i$, so $|E(\mathcal{T})| \leq r^t$.
    
    Now it remains to show that for every $h \in E(\mathcal{H})$ there exists an $h' \in E(\mathcal{T})$ with $h' \subset h$. Observe that at each step $i$, our hyperedge $h$ must intersect $h_i$, so there is a choice of vertex in $h_i \cap h$. Therefore, there is at least one $r$-set constructed that is contained in $h$.
\end{proof}

We first consider the case of minimum positive co-degree at least $2$.

\begin{prop}
Fix $r \geq 3$ and let $n \geq \frac{1}{3} r^4$.
	Let $\mathcal{H}$ be an intersecting $r$-uniform $n$-vertex hypergraph with minimum positive co-degree $\delta^+_{r-1}(\mathcal{H}) \geq 2$. If $\mathcal{H}$ has the maximum number of hyperedges, then $\mathcal{H}$ is a $2$-kernel system.  
\end{prop}

\begin{proof}
	We distinguish three cases based on the minimum transversal size $\tau(\mathcal{H})$ of $\mathcal{H}$.
	
	\smallskip
	
	\noindent {\bf Case 1:} $\tau(\mathcal{H}) = 1$. 
	
	\smallskip
	Then there is a vertex $x$ in each hyperedge of $\mathcal{H}$. Fix a hyperedge $h \in E(\mathcal{H})$ and observe that the $(r-1)$-set $h \setminus \{x\}$ is contained in exactly one hyperedge which violates the positive co-degree condition.
	
	\smallskip
	
	\noindent {\bf Case 2:} $\tau(\mathcal{H}) \geq 3$. 
	
	\smallskip
	
	Then Lemma~\ref{tau-bound} gives
	\[
	|E(\mathcal{H})| \leq r^3 \binom{n-3}{r-3}
	\]
	which for $n \geq \frac{1}{3} r^4$ is smaller than $3 \binom{n-3}{r-2}$, a contradiction.
	\smallskip
	
	\noindent {\bf Case 3:} $\tau(\mathcal{H}) = 2$. 
	
	\smallskip
	
	Let $\{x,y\}$ be a minimum transversal of $\mathcal{H}$. Consider the $(r-1)$-uniform hypergraphs $\mathcal{H}_x = \{h \setminus \{x\} : h \in E(\h) \text{ and }  h \cap \{x,y\} = \{x\}\}$ and $\mathcal{H}_y = \{h \setminus \{y\} : h \in E(\h) \text{ and } h \cap \{x,y\} = \{y\}\}$. First observe that this pair of hypergraphs is cross-intersecting as $\mathcal{H}$ is intersecting. Now observe that any hyperedge $h \in E(\mathcal{H}_x)$ is a set of size $r-1$ that is contained in a hyperedge of $\mathcal{H}$. Thus, $h$ has co-degree at least $2$, therefore must be a member of $\mathcal{H}_y$. This implies that $\mathcal{H}_x = \mathcal{H}_y$, therefore $\mathcal{H}_x$ is intersecting. 
	
	Now if $\h_x = \h_y$ is not a maximal star, then by the Erd\H os-Ko-Rado theorem we have
	\[
	|E(\h)| < 2 \binom{n-3}{r-2} + \binom{n-2}{r-2} = 3 \binom{n-3}{r-2} + \binom{n-3}{r-3},
	\]
	i.e., $\h$ has fewer hyperedges than a $2$-kernel system, a contradiction.
	 Therefore, every hyperedge of $\mathcal{H}_x$ contains a fixed vertex $z$. This implies that every hyperedge of $\mathcal{H}$ contains at least two of $\{x,y,z\}$, i.e., maximality implies that $\mathcal{H}$ is a $2$-kernel system.
\end{proof}

We now turn to the case when $k = 3$. 
We will need two lemmas. The first is due to Frankl (Proposition~1.4 in~\cite{Fr2}).

\begin{lemma}[Frankl,~\cite{Fr2}]\label{cross}
	Let $\A$ and $\B$ be cross-intersecting hypergraphs on vertex set $[N]$ such that $\A$ is $a$-uniform and $\B$ is $(a+1)$-uniform and intersecting. If $N > 2a+1$, then
	\[
	|\A|+|\B| \leq \binom{N}{a},
	\]
	with equality if and only if either $\B$ is empty and $\A$ has size $\binom{N}{a}$ or 
	both $\A$ and $\B$ are maximal stars containing the same a fixed vertex $q$.
\end{lemma}	


The next lemma gives the size of a minimum transversal for a hypergraph with minimum co-degree at least $k$.


\begin{lemma}\label{tau-is-k}
Fix  $r \geq 3$, $k \geq 2$ and let $n \geq 2{\binom{2k-1}{k}}^{-1}(r-k)r^{k+1}$.
	Let $\mathcal{H}$ be an intersecting $r$-uniform $n$-vertex hypergraph with minimum positive co-degree $\delta_{r-1}^+(\mathcal{H}) \geq k$. If $\mathcal{H}$ has the maximum number of hyperedges, then $\mathcal{H}$ has  transversal number $\tau(\mathcal{H}) = k$.
\end{lemma}

\begin{proof}
	First suppose that $\tau(\h) < k$. As in the proof of Lemma~\ref{sunflower-transversal}, let $X$ be a minimal transversal for $\h$ and consider a hyperedge $h$ that intersects $X$ in exactly one element. Such a hyperedge exists as otherwise $X$ is not minimal. The $(r-1)$-set $h \setminus X$ is contained in at most $k-1$ hyperedges of $\h$; one for each element of $X$. This contradicts the co-degree condition on $\h$.
	
	Now suppose that  $\tau(\h) > k$.
	Lemma~\ref{tau-bound} gives
	$|E(\h)| \leq  r^{k+1}\binom{n-k-1}{r-k-1}$. On the other hand, our construction has at least $\binom{2k-1}{k} \binom{n-2k+1}{r-k}$ hyperedges. Therefore, for $n \geq 2{\binom{2k-1}{k}}^{-1}(r-k)r^{k+1}$ we have a contradiction, thus, $\tau(\h)=k$.
\end{proof}

Finally, we need a technical definition to construct auxiliary hypergraphs from $\mathcal{H}$.

\begin{defn}
	Let $\h$ be an $r$-uniform hypergraph and let $T$ be a fixed set of vertices in $\h$.
	For a subset $S \subset T$ define
	\[
	\h_S^T = \{h \setminus S : h \in E(\h) \text{ and }  h \cap T = S\},
	\]
	i.e., $\h^T_S$ is the $(r-|S|)$-uniform hypergraph constructed by removing $S$ from each hyperedge of $\h$ that intersects $T$ in exactly $S$.	
\end{defn}

For ease of notation we will often denote $\h_S^T$ by $\h_{x_1x_2\dots x_s}^T$ when $S = \{x_1,x_2,\dots, x_s\}$.

\begin{thm}\label{k=3}
Fix $r \geq 3$ and let $n \geq 2 r^5$.
	Let $\mathcal{H}$ be an intersecting $r$-uniform $n$-vertex hypergraph with minimum positive co-degree $\delta^+_{r-1}(\mathcal{H}) \geq 3$. 
	If $\mathcal{H}$ has the maximum number of hyperedges, then $\mathcal{H}$ is a $3$-kernel system.
\end{thm}


\begin{proof}
	By Lemma~\ref{tau-is-k} we may assume the minimum transversal size of $\h$ is $\tau(\h)=3$.
	Let $X=\{x,y,z\}$ be a minimum transversal of $\h$. 

	 Consider the three $(r-1)$-uniform hypergraphs $\h^X_x$, $\h^X_y$ and $\h^X_z$. First observe that any pair of these hypergraphs is cross-intersecting as $\h$ is intersecting. Now observe that any hyperedge $h \in E(\h^X_x)$ is a set of size $r-1$ that is contained in a hyperedge of $\h$, therefore $h$ has co-degree at least $3$. This implies that $h$ is also a member of $\h^X_y$ and $\h^X_z$. Thus, all three hypergraphs $\h^X_x, \h^X_y, \h^X_z$ are the same. Moreover, this implies that $\h^X_x$ is intersecting.
	
	We distinguish three cases based on $\tau(\h^X_x)$.
	
	\smallskip
	
	\noindent {\bf Case 1:} $\tau(\h^X_x) =1$.
		
	\smallskip
	
	Let $u$ be a transversal of $\h^X_x$. Every hyperedge of $\h^X_x,\h^X_y,\h^X_z$ contains $u$, therefore, every hyperedge of $\h$ contains at least two vertices from $\{x,y,z,u\}$. 
	Put $T = X \cup \{u\} = \{x,y,z,u\}$.
	 
	 \begin{claim}
	 	The six hypergraphs $\h^T_{ab}$ for $a,b \in T =  \{x,y,z,u\}$ are equal.
	 \end{claim}
	
	\begin{proof}
		It is enough to show that $E(\h^T_{ab}) \subseteq E(\h^T_{ac})$ for any three vertices $a,b,c \in T$.
		Let $h \in E(\h^T_{ab})$ and consider the $(r-1)$-set $h \cup \{a\}$. By the co-degree condition on $\h$ we have that $h \cup\{a\}$ is contained in at least three hyperedges. Each of these hyperedges includes at least two vertices from $\{x,y,z,u\}$, so $h \cup \{a\}$ is contained in the hyperedge $h \cup \{a,c\}$, i.e., 
		$h \in E(\h^T_{ac})$.
	\end{proof}
	
	Observe that $\h^T_{xy}$ and $\h^T_{zu}$ are cross-intersecting, which implies that $\h^T_{xy}$ is intersecting. 
	Now if $\h^T_{xy}$ is not a maximal star, then by the Erd\H os-Ko-Rado theorem we have
	\[
	|E(\h)| < 6 \binom{n-5}{r-3} + 4 \binom{n-4}{r-3} + \binom{n-4}{r-4} = 10 \binom{n-5}{r-3} + 5 \binom{n-5}{r-4} + \binom{n-5}{r-5},
	\]
	i.e., $\h$ has fewer hyperedges than a $3$-kernel system, a contradiction.
	 Therefore, every hyperedge of $\mathcal{H}_{xy}$ contains a fixed vertex $v$. 
	 As the six hypergraphs $\h^T_{ab}$ for $a,b \in T =  \{x,y,z,u\}$ are equal, we have that every hyperedge of $\mathcal{H}$ contains at least three of $\{x,y,z,u,v\}$, i.e., maximality implies that $\mathcal{H}$ is a $3$-kernel system.
	
	\smallskip
	
	\noindent {\bf Case 2:} $\tau(\h^X_x) = 2$.
	
	\smallskip
	
	Let $u,v$ be a minimal transversal of $\h^X_x$, i.e., every hyperedge of $\h^X_x$ contains at least one of $u,v$. As $\h^X_x=\h^X_y=\h^X_z$, we have that every hyperedge of $\h$ contains at least two vertices from $T = \{x,y,z,u,v\}$.
	Moreover, $\h^T_{xu} = \h^T_{yu} = \h^T_{zu}$ and $\h^T_{xv} = \h^T_{yv} = \h^T_{zv}$ and each of these $(r-2)$-uniform hypergraphs is non-empty (as otherwise $u,v$ would not be a minimal transversal).
	Note that there is no hyperedge that intersects $T$ in exactly $u$ and $v$, so $\h^T_{uv}$ is empty. For simplicity, we consider the empty hypergraph as intersecting.
	
	\begin{claim}
		The hypergraph $\h^T_{ab}$ is intersecting for every $a,b \in T = \{x,y,z,u,v\}$.
	\end{claim}
	
	\begin{proof}
		Suppose not. Then there are hyperedges $A,B \in E(\h^T_{ab})$ such that $A \cap B = \emptyset$.
		By the co-degree condition, the $(r-1)$-set $A \cup \{a\}$ is contained in at least three hyperedges of $\h$. Since each hyperedge of $\h$ contains at least two elements from $T$, there is a hyperedge $A \cup \{a, c\}$ where $c \in T\setminus \{a,b\}$. Similarly, the $(r-1)$-set $B \cup \{b\}$ is contained in some hyperedge $B \cup \{b , d\}$  where $d \in T \setminus \{a,b,c\}$.
		However, the hyperedges $A \cup \{a,  c\}$ and $B \cup \{b, d\}$ are disjoint which violates the intersecting property of $\h$.
	\end{proof}
	
		Now for any $a,b \in T$ we have  $\h^T_{T\setminus\{a,b\}}$ and $\h^T_{ab}$ are cross-intersecting, $\h^T_{T\setminus\{a,b\}}$ is $(r-3)$-uniform and $\h^T_{ab}$ is $(r-2)$-uniform and intersecting. Therefore, as $n-5 > 2(r-3)+1$, 
		we may apply Lemma~\ref{cross} to get
	\[
	|E(\h^T_{ab})| + |E(\h^T_{T\setminus\{a,b\}})| \leq \binom{n-5}{r-3}.
	\]
Thus
\[
|E(\h)| = \sum_{S\subseteq T} |E(\h^T_S)|  \leq 10 \binom{n-5}{r-3} + 5 \binom{n-5}{r-4} + \binom{n-5}{r-5}.
\]

As $\h$ has the maximum number of hyperedges, we must have equality above.
Therefore, we must have that for every $a,b \in T$, the hypergraphs $\h^T_{T\setminus\{a,b\}}$ and $\h^T_{ab}$ have the form of one of the two extremal constructions in Lemma~\ref{cross}. In particular, $\h^T_{ab}$ is either empty or  a maximal star. As $\h^T_{xu} = \h^T_{yu} = \h^T_{zu}$ and $\h^T_{xv} = \h^T_{yv} = \h^T_{zv}$ are non-empty, each is a maximal star. The hypergraphs $\h^T_{xu}$ and $\h^T_{yv}$ are cross-intersecting which implies that all six of these these maximal stars share the same fixed vertex $q$. Therefore, we can replace minimal transversal $u,v$ of $\h_x^X$ with $q$, a contradiction.


	
	\smallskip

	\noindent {\bf Case 3:} $\tau(\h^X_x) \geq 3$.
	
	\smallskip
	
	Then Lemma~\ref{tau-bound} gives
	\[
	|E(\h^X_x)| \leq (r-1)^3 \binom{(n-1)-3}{(r-1)-3} \leq r^3\binom{n-4}{r-4}.
	\]
	The remaining hyperedges of $\h$ are counted by $\h^X_{xyz}$ and $\h^X_{ab}$ for $a,b \in \{x,y,z\}$. We need a simple claim. Recall that the {\it shadow} of an $r$-uniform hypergraph $\mathcal{G}$ is the collection of all $(r-1)$-sets contained in a hyperedge of $\mathcal{G}$. We denote the shadow of $\mathcal{G}$ by $\Delta(\mathcal{G})$.
	
	\begin{claim}\label{shadow}
		For each hyperedge $h \in E(\h^X_{yz})$ there is some hyperedge $g \in E(\h^X_x)$ that contains $h$. Thus,
		\[
		|E(\h^X_{yz})| \leq |\Delta(\h^X_x)|.
		\]
	\end{claim}

	\begin{proof}
	    Let $h$ be an arbitrary hyperedge of $\h^X_{yz}$.
		Consider the $(r-1)$-set $A = h \cup \{y\}$. The set $A$ has co-degree at least $3$, so it is contained in three hyperedges of $\h$; one such hyperedge is $A \cup \{z\}$, another could be $A \cup \{x\}$, so there exists at least one hyperedge of the form $A \cup \{w\}$ where $w \not \in \{x,y,z\}$. However, $A \cap \{x,y,z\} = \{y\}$, so $(A \cup \{w\}) \setminus \{y\} \in E(\h^X_y) = E(\h^X_x)$.
	\end{proof}
	
	By Claim~\ref{shadow} we have
	\[
	|E(\h^X_{yz})| \leq |\Delta(\h^X_x)| \leq (r-1) |E(\h^X_x)| \leq r^4\binom{n-4}{r-4}.
	\]
	Finally, $|E(\h^X_{xyz})| \leq \binom{n-3}{r-3}$. Thus,
	\[
	|E(\h)| \leq \binom{n-3}{r-3} + 3(r^4+r^3)\binom{n-4}{r-4}
	\]
	which is less than $10 \binom{n-5}{r-3}$ for $n \geq 2r^5$, a contradiction.
\end{proof}

In order to extend the technique used in this section to reprove our theorem for minimum positive co-degree $k \geq 4$ we would need to distinguish additional cases based on the transversal size of $\h_x^X$. Some of these cases can be addressed with Lemmas~\ref{tau-bound} and \ref{cross}, but probably new ideas will be needed.


\section*{Acknowledgements}
The authors would like to thank Dhruv Mubayi for pointing out reference~\cite{MZ2} and Lemma~\ref{k-core}. We also thank the anonymous referees for their careful reading of the manuscript and many helpful comments that improved the presentation.

\section*{Appendix}

We now give an improvement to Lemma~\ref{k-core} which we believe is of independent interest. Recall that $f(r,p)$ is the minimum integer such that an $r$-uniform hypergraph with $f(r,p)$ hyperedges contains a sunflower with $p$ petals.

\begin{lemma}\label{k-core-new}
Fix integers $r\geq 3$, $k \geq 1$ and $p \geq 1$ and let $n$ be large enough.
If $\mathcal{G}$ is an $r$-uniform $n$-vertex hypergraph with 
\[
|E(\mathcal{G})| \geq 2 r^{r-k} f(r,p r^{r-k}) \binom{n-k-1}{r-k-1},
\]
then  $\mathcal{G}$ contains a sunflower with $p$ petals and core of size at most $k$.  
\end{lemma}


This replaces the value of $C=(pr2^r)^{2^r}$ in Lemma~\ref{k-core} with $C=2 r^{r-k} f(r,p r^{r-k})$ which is significantly smaller when using the bound on $f(r,pr^{r-k})$ from \cite{ALWZ}.

\begin{proof}
For the sake of a contradiction, suppose that $\mathcal{G}$ contains no sunflower with $p$ petals and core of size at most $k$.

Iteratively remove from $\mathcal{G}$ a sunflower $\mathcal{S}$ with exactly $p  r^{c(\mathcal{S})-k}$ petals such that at each step we choose a sunflower with minimum available core size $c(\mathcal{S})$. Let $t$ be the number of steps in this sunflower removal procedure. Note that $t$ grows with $n$ as at each step we remove at most $p r^{r-k}$ hyperedges from $\mathcal{G}$ and we only need constant number of hyperedges to guarantee the existence of a sunflower with $p  r^{c(\mathcal{S})-k}$ petals. In particular, we have
\[
t \geq \frac{|E(\mathcal{G})|-f(r,p  r^{r-k})}{p  r^{r-k}} \geq \frac{|E(\mathcal{G})|}{2pr^{r-k}}
\]
for $n$ large enough.

The core of each removed sunflower is of size at least $k+1$ and at most $r-1$. Therefore, there is some integer $s$ such that there are at least $t /r$ cores of size $s$ among the removed sunflowers.
 Some of these cores may be identical. Let us compute the maximum multiplicity of a core $Y$. There are at most $\binom{n-|Y|}{r-|Y|}$ hyperedges containing $Y$ and each removed sunflower with core $Y$ has exactly $p r^{|Y|-k}$ hyperedges. Therefore,
 the maximum multiplicity of a core $Y$ is at most
\[
\frac{1}{p  r^{|Y|-k}}\binom{n-|Y|}{r-|Y|} \leq \frac{1}{p  r} \binom{n-k-1}{r-k-1}
\]
for $n\geq r$.
Therefore, there is a collection of at least 
\[
(t/r)  p  r \binom{n-k-1}{r-k-1}^{-1} \geq p  \frac{|E(\mathcal{G})|}{2pr^{r-k}}  \binom{n-k-1}{r-k-1}^{-1} \geq f(r,p  r^{r-k})
\]
distinct cores of size $s$. Let $Y_1,Y_2,\dots, Y_q$ be these cores and let $\mathcal{S}_i$ be the sunflower with  core $Y_i$ for $i=1,2,\dots,q$. Note that each of these sunflowers has exactly $p  r ^{s-k}$ petals.

Let $t$ be the first step in the sunflower removal procedure in which a sunflower with core of size $s$ is chosen to be removed. This implies that all later cores are of size at least $s$. Now we will show that there is a sunflower $\mathcal{B}$ with core of size less than $s$ and $p r^{c(\mathcal{B})-k}$ petals among the hyperedges in the sunflowers $\mathcal{S}_1,\mathcal{S}_2,\dots, \mathcal{S}_q$. Before removing the sunflower in step $t$, all hyperedges of the sunflowers $\mathcal{S}_1,\mathcal{S}_2,\dots, \mathcal{S}_q$ are still in $\mathcal{H}$. Therefore, the sunflower $\mathcal{B}$ with core of size less than $s$ could be chosen in step $t$, this will contradict the choice of $t$.

We may think of the $s$-sets $Y_1,\dots, Y_q$ as an $s$-uniform hypergraph on the vertex set of $\mathcal{H}$.
As $q \geq f(r,p  r^{r-k})\geq f(s,p  r^{r-k}) \geq f(s,pr^{s-k})$, the $s$-sets $Y_1,\dots, Y_q$ contain an $s$-uniform sunflower $\mathcal{A}$ with $p  r^{s-k}$ petals and core $Y^*$ of size less than $s$. By relabelling, we may suppose that $Y_i$ is a member of $\mathcal{A}$ for $i=1,2,\dots, p  r^{s-r}$. Note that the petals $Y_i \setminus Y^*$ of $\mathcal{A}$ are pairwise disjoint by definition. The sunflower $\mathcal{A}$ is not in the hypergraph $\mathcal{H}$ as it is $s$-uniform. However, each hyperedge of $\mathcal{A}$ is the core of some sunflower $\mathcal{S}_i$ in $\mathcal{H}$. Therefore, we will use the members of $\mathcal{A}$ to identify an $r$-uniform sunflower $\mathcal{B}$ with core $Y^*$ in $\mathcal{H}$. The main idea will be carefully choose a petal from each sunflower $\mathcal{S}_i$ whose core is a member of $\mathcal{A}$. To this end, define $\mathcal{B}$ as follows:

First pick any hyperedge of $\mathcal{S}_1$; denote it by $h_1$. Now suppose we have chosen $\ell$ hyperedges $h_1,h_2,\dots, h_\ell$ that form a sunflower with core $Y^*$. The union of these hyperedges contains $\ell(r-|Y^*|)$ vertices outside of $Y^*$. Therefore, as long as 
\begin{equation}\label{ell-bound}
p  r ^{s-k}> \ell(r-|Y^*|),
\end{equation}
there is a petal $Y_i \setminus Y^*$ of $\mathcal{A}$ that is disjoint from each of the hyperedges $h_1,h_2,\dots, h_\ell$. The corresponding sunflower $\mathcal{S}_i$ with core $Y_i$ has 
\[
p  r ^{s-k} > \ell ({r-|Y^*|})
\]
petals by (\ref{ell-bound}). Therefore, there is a petal $P$ of $\mathcal{S}_i$ that is also disjoint from the hyperedges in $h_1,h_2,\dots, h_\ell$. Let $h_{\ell+1}$ be the hyperedge $P \cup Y_i$. Now we have a sunflower with $\ell+1$ petals and core $Y^*$. We may repeat this procedure as long as $\ell$ satisfies (\ref{ell-bound}), i.e., until $\ell = p  r^{s-k-1}$ . This implies that the number of petals in sunflower $\mathcal{B}$ is at least 
\[
p  r^{s-k-1}.
\]
As $\mathcal{B}$ has core $Y^*$ of size $c(\mathcal{B}) < s$ we have a contradiction to the choice of sunflower in step $t$.
\end{proof}

\end{document}